OpenStudy (anonymous):

Given f(x) = 3x and g(x) = x – 7, (f g)(x) =

OpenStudy (anonymous):

helppppp

OpenStudy (mimi_x3):

\[f(x) = 3x ; g(x) = x-7\] \(f(g(x)) = f(x-7) = 3(x-7) = 3x-21\)

OpenStudy (mimi_x3):

@beezz12: you get it?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@beezz12 what help its done